OpenStudy (anonymous):

How do you find the range of a function? I know how to find the domain, but how do you find the range of a function algebraically without a graph? Please don't give me an easy example and solve it. Instead what method should I use to get a reliable answer every time?

OpenStudy (anonymous):

you can find horizontal asymptotes, determine what happens as x approaches infinity and negative infinity

OpenStudy (anonymous):

when you take calculus you'll be able to find absolute max and absolute min

OpenStudy (anonymous):

\[f(x) = \sqrt{-x^2 - x + 6}\] How do you find the range of that? I know the domain is x is not -3, 2

OpenStudy (anonymous):

that's not the correct domain

OpenStudy (anonymous):

you have the critical numbers though, which is a start

OpenStudy (anonymous):

you found the zeros of -x^2-x+6

OpenStudy (anonymous):

but -x^2-x+6 can't be negative

OpenStudy (anonymous):

do you know how to test for that?

OpenStudy (anonymous):

I don't

OpenStudy (anonymous):

ok the zeros you found divide the x-axis into intervals

OpenStudy (anonymous):

they are the x-intercepts obviously

OpenStudy (anonymous):

so the intervals you are to consider are these:

OpenStudy (anonymous):

\[(-\infty, -3), (-3, 2), (2, \infty)\]

OpenStudy (anonymous):

Ok. How do you find the domain from that though?

OpenStudy (anonymous):

now we pick a random number from these intervals and check the sign of f(x) in that interval. all numbers in the interval will have the same sign

OpenStudy (anonymous):

we can reconstruct the polynomial using the zeros you found

OpenStudy (anonymous):

\[(x+3)(x-2)\ge 0\]

OpenStudy (anonymous):

so now lets see what happens when we pick a value from the interval (-infinity, -3)

OpenStudy (anonymous):

i'll use -4 for this example

OpenStudy (anonymous):

(-4+3)(-4-2)= a positive number, so its greater than 0

OpenStudy (anonymous):

negative*negative=positive

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

right, so now you know how to do that, we can use a number line

OpenStudy (anonymous):